# Extreme time extrapolation capabilities and thermodynamic consistency of physics-inspired Neural Networks for the 3D microstructure evolution of materials

### v1

@@ Abstract
A Convolutional Recurrent Neural Network (CRNN) is trained to reproduce the evolution of the spinodal decomposition process in three dimensions as described by the Cahn-Hilliard equation. A specialized, physics-inspired architecture is proven to provide close accordance between the predicted evolutions and the ground truth ones obtained via conventional integration schemes. The method can closely reproduce the evolution of microstructures not represented in the training set at a fraction of the computational costs. Extremely long-time extrapolation capabilities are achieved, up to reaching the theoretically expected equilibrium state of the system, despite the training set containing only relatively-short, initial phases of the evolution. Quantitative accordance with the decay rate of the Free energy is also demonstrated up to late coarsening stages, providing an example of a data-driven, physically consistent and high-accuracy Machine Learning method for the long timescale simulation of materials.
preprint: AIP/123-QED
@@ I Introduction
Recently, there has been a surge in interest in Machine Learning (ML) methods bishop2006pattern ; goodfellowdeep2016 in the Computational Physics and Materials Science community butler2018machine ; Mehta2019PhysRep ; bedolla2020machine ; NGUYEN2024100544 . For instance, the new field of ML interatomic potentials Bartok2010PRL ; Kocer2022review promises to push the boundaries of tractable time and spatial scales in molecular dynamics. An initially less explored route, which has however gained traction Kim2019CGF ; Fulton2019CGF ; Zhang2020CMAME ; MontesdeOcaZapiain2021npj ; Yang2021Patterns ; Lanzoni2022PRM ; LUIS , stems in the possibility of leveraging ML approaches, and Neural Networks (NN) in particular, for meso- and macroscopic models. Indeed, while it is easier to reach experimental scales using such tools, computational costs might still be critical for stiff problems and fine discretization requirements. ML may therefore provide an alternative to conventional solvers and allow one to bypass these issues. Indeed, in the last years, efforts have been dedicated to the possibility of approximating Partial Differential Equations (PDEs) with Machine Learning models RAISSI2019PINN ; Bretin2022JCompPhys , with a particular interest on the morphological and microstructural evolution of materials Yang2021Patterns ; peivaste2022machine ; Lanzoni2022PRM ; fan2024accelerate .
@@
In this work, we explore the possibility of using a Convolutional Recurrent Neural Network (CRNN) to reproduce the time-dependent solutions of the Cahn-Hilliard equation, which describes the spinodal decomposition, an important process in binary mixtures leading to a spontaneous separation into two different phases Langer1971AnnPhy ; Kwon2007PRE ; Andrews2020PRM ; jinnai2000geometrical , focusing on their long-time behavior. As the concentration fields can be conveniently mapped into a single order parameter, $\varphi$, the considered model also provides a useful testing ground for a more general class of Phase Field (PF) models Li2009CommunComput ; provatas2011phase , which are particularly interesting for their flexibility in the description of materials evolution. This opens the way for future application to multi-physics approaches involving several contributions to the driving force of mesoscale behaviors.
@@
A 2D version of the spinodal decomposition has already been addressed via ML methods by the community in previous works, also using CRNNs MontesdeOcaZapiain2021npj ; Yang2021Patterns . However, tackling the full 3D problem is essential to model realistic materials. Steps in this direction can be found in a few, very recent publications. In Ref. wang2023machine, a Convolutional (but not Recurrent) NN is shown to nicely predict relatively few steps of spinodal decomposition evolution, sufficient for the specific application therein. Ref. fan2024accelerate, , instead, shows that a Recurrent, Graph Network can surrogate grain coarsening dynamics in 2D and 3D. In the same work, an adaptive time-stepping scheme is proposed, which allows one to obtain an even larger speed-up in predictions. Still, only sequences of some hundred states are shown for the 3D case. In the present work, we therefore focus on the possibility of obtaining a model capable of producing stable, extremely long microstructural evolutions that are consistent with the underlying thermodynamic assumptions. This is done by exploiting a physics-inspired layer, which closely mimics the underlying material flow process in the ML model architecture. Importantly, predicted sequences are shown to reach the correct stationary state of the system, despite the CRNN only being presented short, far from equilibrium sequences during training.
@@ II.1 Phase Field model and dataset generation
Spinodal decomposition consists in the spontaneous separation of a binary mixture in two different phases below a critical temperature. From a mathematical standpoint, it is convenient to map the concentration fields tracking the local chemical composition to a single scalar field $\varphi$. This way, the dynamics of the system can be described in a diffuse interface, PF model provatas2011phase . The temporal evolution of the system may be derived from the definition of the free energy functional $F[\varphi]$:
@@
where $\Omega$ represents the physical domain considered, $g(\varphi)$ is the "bulk" contribution to the free energy, and the gradient term $|\vec{\nabla}\varphi|^{2}$ is related to interface energy and width $\varepsilon$ Li2009CommunComput ; provatas2011phase . For the bulk term $g$, we choose the common double-well potential expression $g(\varphi)=\frac{18}{\varepsilon}\varphi^{2}(1-\varphi)^{2}$ Li2009CommunComput .
The equations of motion may be derived from a variational principle provatas2011phase . $\varphi$ evolves in the direction of free-energy minimization, resulting in the well-known Cahn-Hilliard equation cahn1958free ; cahn1965JCP :
@@
where $g^{\prime}=dg/d\varphi$, $M$ is a mobility constant and $\vec{J}$ is $\varphi$ diffusion flux, proportional to the gradient of the generalized chemical potential $\mu=\delta F/\delta\varphi$. Among PF models, the Cahn-Hilliard equation is particularly well suited as a relatively simple testing ground for advanced ML methods, as it can exhibit rich, non-linear behavior during evolution despite the relatively simple mathematical formulation, e.g. resulting in the formation and subsequent coarsening of 3D patterns, which are also interesting in the context of metamaterials design kumar2020inverse ; wang2023machine .
The above-defined model is herein exploited to create a dataset of sequences of microstructure evolutions. Eq. 2 is solved with a finite difference scheme on a cubic $64\times 64\times 64$ grid using an explicit forward-Euler method with a constant integration step of $\delta t=1.25\times 10^{-3}$. $M$ is set to unity and $\varepsilon$ to three grid units. These parameters are chosen because they offer a good compromise between the accuracy of the solution of the PDE and the time required to generate the dataset. Periodic boundary conditions (PBCs) are imposed. As starting configurations for the field, we consider periodic Perlin noise samples Perlin1985287 , obtained by suitable adaptation of the Python project at Ref. PythonPerlinNoise, . Perlin noise is a type of gradient noise that allows one to effectively select the typical correlation length of features in random $\varphi$ states. This is in contrast with the simpler initialization with uniform, random $\varphi$ values typically used, but provides coarser, grid-independent initial conditions that are easier for the CRNN to correlate with subsequent states of the evolution.
The generated training set is composed of 1850 sequences with different average $\varphi$ and correlation lengths, which are then split into a 70:30 ratio, forming the training and validation sets. The resulting average value of $\varphi$, $\overline{\varphi}$, has a bell-shaped distribution with mean $1/2$ and a standard deviation of $\approx 1/10$ as a result of the different Perlin noise parameters, ensuring the variability of initial conditions. Each sequence is composed of 50 consecutive configurations separated by a time interval $\tau=0.5$ in time units (corresponding to 400 $\delta t$ integration steps). Training sequences start either from the initial Perlin noise or later states in the evolution, allowing the NN to generalize to partially-coarsened microstructures as initial conditions.
@@ II.2 NN structure and training
Figure 1: (a) Sketch of the NN architecture and specification of the "toFlux" layer converting the hidden state $h^{2}_{t+\tau}$ to the vector field $\vec{u}_{t}$. (b) Lossplot for training the Convolutional Recurrent Neural Network. No sign of overfitting is present. Differences in initial values for the curves are expected and due to the training procedure (see text for more information). (c) Comparison between the ground truth evolution and the predicted one for a validation set case. In the upper half of the snapshots, only $\varphi\geq 0.5$ is shown to better appreciate the internal structure.
The core architecture of the CRNN used in this work is based on Ref. Lanzoni2022PRM, but for three critical modifications. First, since we are considering 3D evolutions in the present work, all convolutional layers are promoted to their three-dimensional counterparts, using the corresponding PyTorch paszkepytorch2019 implementation. Circular padding schubert2019circular enforces PBCs, but alternative boundary conditions could be straightforwardly translated into other padding modes. Second, a deeper mapping between the input, hidden and output layers in the recurrent modules of the CRNN (see GitHub repository for technical details) is exploited to increase the representation capabilities of the network. Third, a new output layer inspired by the physical formulation of the underlying problem is introduced to better account for the material flow dynamics of Eq. 2. More precisely, we use a simple forward-Euler integration scheme ansatz and, instead of directly predicting the next state of the field $\hat{\varphi}_{t+\tau}$ (hat denotes predicted quantities), we define the NN output as the vector field $\vec{u}_{t}$, such that:
@@
Here $\vec{\nabla}\cdot\vec{u}_{t}$ is calculated using a finite difference scheme, efficiently implemented as non-trainable convolutional layers with circular padding and zero bias. Moreover, due to PBCs, $\vec{J}$ should have a vanishing mean, which in the NN architecture is translated as removing from $\vec{u}_{t}$ its mean value. We remark that this approach, while resembling a forward-Euler integration scheme, is strictly more powerful, as the NN can exploit recurrence to access information farther in the evolution past, thus being more similar to a high-order integration scheme. More importantly, this procedure ensures the exact conservation of $\varphi$ by construction, thus making it suitable for other diffusive processes, such as the surface diffusion case of Ref. Lanzoni2022PRM, , where $\varphi$ conservation was enforced only weakly through the loss function. A similar strategy using an intermediate vector field was also exploited at Ref. Kim2019CGF, for incompressible flows.
@@
This physics-inspired layer is in line with the well-known and used fact that good inductive biases encoded in the NN architecture lead to an easier training procedure, better generalization capabilities and a lower number of parameters, which in turn reduces overfitting risks goodfellowdeep2016 ; Mehta2019PhysRep . The present choice of fully convolutional layers long2015fully itself serves as an encoding of the translational equivariance symmetry Cohen2016groupEquivariant and of the locality of the evolution, allowing for applications to arbitrary domain sizes.
@@
Based on our tests, the best set of hyperparameters is the following: the CRNN is composed of two, stacked Convolutional Gated Recurrent Unit (GRU) Chung20141arXiv ; Shi2015NIPS blocks, each using 10 channels for hidden states and $5\times 5\times 5$ convolution kernels. A sketch of the overall CRNN architecture is shown in Fig. 1(a). The resulting NN model contains $\approx 2.9\times 10^{4}$ trainable parameters, almost an order of magnitude less than the 2D prototype presented in Ref. Lanzoni2022PRM, , despite the more challenging 3D setting here tackled. Such compression is made possible by the modifications previously discussed. The training loss used is the voxel-wise Mean Squared Error between the predicted and ground truth sequence. Training of the NN parameters is performed using the standard implementation of the Adam optimizer kingmaadam2014 with a learning rate of $10^{-5}$. Before being passed to the CRNN, training sequences are downscaled by a factor of $1/2$ in all spatial dimensions using nearest interpolation to ease GPU memory requirements during training. As it will be shown in the following, the NN prediction accuracy is not significantly impacted by such a downscale process, which, on the other hand, can be leveraged to tackle larger computational cells at reduced computational costs, as has also been demonstrated in previous works Lanzoni2022PRM ; fan2024accelerate . Data augmentation based on reflections, 90°90°90  ° degree rotations and the $\mathbf{Z}_{2}$ symmetry $\varphi\rightarrow(1-\varphi)$ is also exploited in training. We point again readers interested in implementation details to the GitHub repository, where the full code used to perform training and evolutions here reported is available.
@@
Fig. 1(b) reports training and validation loss respectively. Discrepancies between the two values in the initial training stages are due to the use of a curriculum learning schedule bengio2009curriculum ; goodfellowdeep2016 , which proved to speed up training convergence. Specifically, at the beginning of training the model parameters are optimized using a simplified loss function, i.e. the CRNN is required to generate only the last state provided the previous 49 ones. In the following 25 epochs, the loss function gets gradually more complex as the CRNN is required to generate progressively longer sequences. At the end of this curriculum stage, the ML model has to reproduce the full sequence based only on the initial condition.
@@
Figure 2: (a) Spatial generalization test. The domain size is 8 times larger than that used in the training set (twice as large in all directions). One-to-one accordance with the ground truth evolution can still be observed. (b) Evolution on a computational cell with the same size of training for a $\varphi$ initial profile shaped as the Stanford bunny. In the upper half of the snapshots, only $\varphi\geq 0.5$ is shown.

### v2

@@ Abstract
A Convolutional Recurrent Neural Network (CRNN) is trained to reproduce the evolution of the spinodal decomposition process in three dimensions as described by the Cahn-Hilliard equation. A specialized, physics-inspired architecture is proven to provide close accordance between the predicted evolutions and the ground truth ones obtained via conventional integration schemes. The method can accurately reproduce the evolution of microstructures not represented in the training set at a fraction of the computational costs. Extremely long-time extrapolation capabilities are achieved, up to reaching the theoretically expected equilibrium state of the system, consisting of a layered, phase-separated morphology, despite the training set containing only relatively-short, initial phases of the evolution. Quantitative accordance with the decay rate of the Free energy is also demonstrated up to the late coarsening stages, proving that this class of Machine Learning approaches can become a new and powerful tool for the long timescale and high throughput simulation of materials, while retaining thermodynamic consistency and high-accuracy.
@@ I Introduction
Recently, there has been a surge in interest in Machine Learning (ML) methods Bishop (2006); Goodfellow et al. (2016) in the Computational Physics and Materials Science community Butler et al. (2018); Mehta et al. (2019); Bedolla et al. (2020); Nguyen et al. (2024). For instance, the new field of ML interatomic potentials Bartók et al. (2010); Kocer et al. (2022) promises to push the boundaries of tractable time and spatial scales in molecular dynamics. An initially less explored route, which has however gained traction Kim et al. (2019); Fulton et al. (2019); Zhang and Garikipati (2020); Montes de Oca Zapiain et al. (2021); Yang et al. (2021); Lanzoni et al. (2022); Martín-Encinar et al. (2024), stems in the possibility of leveraging ML approaches, and Neural Networks (NN) in particular, for meso- and macroscopic models. Indeed, while it is easier to reach experimental scales using such tools, computational costs might still be critical for stiff problems and fine discretization requirements. ML may therefore provide an alternative to conventional solvers and allow one to bypass these issues. Indeed, in the last years, efforts have been dedicated to the possibility of approximating Partial Differential Equations (PDEs) with Machine Learning models Raissi et al. (2019); Bretin et al. (2022), with a particular interest on the morphological and microstructural evolution of materials Yang et al. (2021); Peivaste et al. (2022); Lanzoni et al. (2022); Fan et al. (2024).
@@
In this work, we explore the possibility of using a Convolutional Recurrent Neural Network (CRNN) to reproduce the time-dependent solutions of the Cahn-Hilliard equation, which describes the spinodal decomposition, an important process in binary mixtures leading to a spontaneous separation into two different phases Langer (1971); Kwon et al. (2007); Andrews et al. (2020); Jinnai et al. (2000), focusing on their long-time behavior. As the concentration fields can be conveniently mapped into a single order parameter, $\varphi$, the considered model also provides an initial test for the more general class of Phase Field (PF) models Li et al. (2009); Provatas and Elder (2011). This is advantageous as PF models proved to be effective in several fields of Materials Science Chen (2002); Boettinger et al. (2002); Li et al. (2009); Provatas and Elder (2011); Albani et al. (2016), thanks to their natural ability to tackle complex geometries, eventually involving topological changes, such as domain coalescence and splitting. Additionally, their formulation is open to the addition of multiple physical contributions, thus making the development of NN surrogates particularly appealing.
@@
A 2D version of the spinodal decomposition has already been addressed via ML methods by the community in previous works, also using CRNNs Montes de Oca Zapiain et al. (2021); Yang et al. (2021). However, tackling the full 3D problem is essential to model realistic materials. For instance, the 3D coarsening of the spinodal decomposition may exhibit bi-continuous patterns, under suitable conditions, which cannot be observed in 2D. Additionally, full 3D simulations are computationally more challenging due to the requirement of a higher number of collocation points in discretization procedures, which further calls for acceleration with novel tools. Steps in this direction can be found in a few, very recent publications. In Ref. Wang et al., 2023 a Convolutional (but not Recurrent) NN is shown to nicely predict relatively few steps of spinodal decomposition evolution, sufficient for the specific application therein. Ref. Fan et al., 2024, instead, shows that a Recurrent, Graph Network can surrogate grain coarsening dynamics in 2D and 3D. In the same work, an adaptive time-stepping scheme is proposed, which allows one to obtain an even larger speed-up in predictions. Still, only sequences of some hundred states are shown for the 3D case. In the present work, we therefore focus on the possibility of obtaining a model capable of producing stable, extremely long microstructural evolutions that are consistent with the underlying thermodynamic assumptions. This is done by exploiting a physics-inspired layer, which closely mimics the underlying material flow process in the ML model architecture. Importantly, predicted sequences are shown to reach the correct stationary state of the system, despite the CRNN only being presented orders of magnitude shorter, far from equilibrium sequences during training.
@@ II.1 Phase Field model and dataset generation
Spinodal decomposition consists in the spontaneous separation of a binary mixture in two different phases below a critical temperature. From a mathematical standpoint, it is convenient to map the concentration fields tracking the local chemical composition to a single scalar field $\varphi$. This way, the dynamics of the system can be described in a diffuse interface, PF model Provatas and Elder (2011). The temporal evolution of the system may be derived from the definition of the free energy functional $F[\varphi]$:
@@
where $\Omega$ represents the physical domain considered, $g(\varphi)$ is the "bulk" contribution to the free energy, and the gradient term $|\vec{\nabla}\varphi|^{2}$ is related to interface energy and width $\varepsilon$ Li et al. (2009); Provatas and Elder (2011). As we are not interested in modeling a specific material, we choose for the bulk term $g$ the common and numerically convenient double-well potential expression $g(\varphi)=\frac{18}{\varepsilon}\varphi^{2}(1-\varphi)^{2}$ Li et al. (2009).
The equations of motion may be derived from a variational principle Provatas and Elder (2011). $\varphi$ flows in the direction of free-energy minimization, resulting in the well-known Cahn-Hilliard equation Cahn and Hilliard (1958); Cahn (1965):
@@
where $g^{\prime}=dg/d\varphi$, $M$ is a mobility constant and $\vec{J}$ is $\varphi$ diffusion flux, proportional to the gradient of the generalized chemical potential $\mu=\delta F/\delta\varphi$. Among PF models, the Cahn-Hilliard equation is particularly well suited as a relatively simple testing ground for advanced ML methods, as it can exhibit rich, non-linear behavior during evolution despite the relatively simple mathematical formulation, e.g. resulting in the formation and subsequent coarsening of 3D patterns, which are also interesting in the context of metamaterials design Kumar et al. (2020); Wang et al. (2023).
The above-defined model is herein exploited to create a dataset of sequences of microstructure evolutions. Eq. 2 is solved with a finite difference scheme on a cubic $64\times 64\times 64$ uniform grid using an explicit forward-Euler method with a constant integration step of $\delta t=1.25\times 10^{-3}$. $M$ is set to unity and $\varepsilon$ to 3 grid units. These parameters are chosen because they offer a good compromise between the accuracy of the solution of the PDE and the time required to generate the dataset. Periodic boundary conditions (PBCs) are imposed. As starting configurations for the field, we consider periodic Perlin noise samples Perlin (1985), obtained by suitable adaptation of the Python project at Ref. Pyt, 2023. Perlin noise is a type of gradient noise that allows one to effectively select the typical correlation length of features in random $\varphi$ states. This is in contrast with the simpler initialization with uniform, random $\varphi$ values typically used, but provides coarser, grid-independent initial conditions that are easier for the CRNN to correlate with subsequent states of the evolution. Indeed, pure white noise cannot be recognized by CRNN as a sufficiently featured input, acting as an ill-defined initial condition and leading to arbitrary subsequent morphologies.
The generated training set comprises 1850 sequences with different average $\varphi$ values and correlation lengths, which are then split into a 70:30 ratio, forming the training and validation sets. The resulting average value of $\varphi$, $\overline{\varphi}$, has a bell-shaped distribution with mean $1/2$ and a standard deviation of $\approx 1/10$ as a result of the different Perlin noise parameters, ensuring the variability of initial conditions. Each sequence is composed of 50 consecutive configurations separated by a time interval $\tau=0.5$ in time units (corresponding to 400 $\delta t$ integration steps). Training sequences start either from the initial Perlin noise or later states in the evolution, allowing the NN to generalize to partially coarsened microstructures as initial conditions.
@@ II.2 NN structure and training
Figure 1: (a) Sketch of the NN architecture and specification of the "toFlux" layer converting the hidden state $h^{2}_{t+\tau}$ to the vector field $\vec{u}_{t}$. (b) Lossplot obtained training the Convolutional Recurrent Neural Network. No sign of overfitting is present. Differences in initial values for the curves are expected and due to the training procedure (see text for more information). (c) Comparison between the ground truth evolution and the predicted one for a validation set case. In the upper half of the snapshots, only $\varphi\geq 0.5$ is shown to better appreciate the internal structure.
The core architecture of the CRNN used in this work is based on Ref. Lanzoni et al., 2022 but for three critical modifications. First, since we are considering 3D evolutions in the present work, all convolutional layers are promoted to their three-dimensional counterparts, using the corresponding PyTorch Paszke et al. (2019) implementation. Circular padding Schubert et al. (2019) enforces PBCs, but alternative boundary conditions could be straightforwardly translated into other padding modes. Second, a deeper mapping between the input, hidden and output layers in the recurrent modules of the CRNN (see GitHub repository for technical details) is exploited to increase the representation capabilities of the network. Third, a new output layer inspired by the physical formulation of the underlying problem is introduced. In particular, our main concern here is to strongly constrain the CRNN prediction to be consistent with conservative flow dynamics of Eq.2, a property that standard Convolutional layers do not possess. To this end, we use a simple forward-Euler integration scheme ansatz and, instead of directly predicting the next state of the field $\hat{\varphi}_{t+\tau}$ (hat denotes predicted quantities), we define the NN output as the vector field $\vec{u}_{t}$, such that:
@@
where $\vec{u}_{t}$ is a vector field acting as the flux term $-\tau\vec{J}_{t}$ in the standard forward-Euler integration scheme of Equation 2. Here $\vec{\nabla}\cdot\vec{u}_{t}$ is calculated using a finite difference scheme, efficiently implemented as non-trainable convolutional layers with circular padding and zero bias. Moreover, due to PBCs, $\vec{J}$ should have a vanishing mean, which in the NN architecture is translated as removing from $\vec{u}_{t}$ its mean value. While resembling a forward-Euler integration scheme, we remark that this approach is strictly more powerful, as the NN can exploit recurrence to access information farther in the evolution past, thus being more similar to a high-order integration scheme. More importantly, this procedure ensures the exact conservation of $\varphi$ by construction, thus making it suitable for other diffusive processes, such as the surface diffusion case of Ref. Lanzoni et al., 2022, where $\varphi$ conservation was enforced only weakly through the loss function. A similar strategy using an intermediate vector field was also exploited at Ref. Kim et al., 2019 for incompressible flows.
@@
This physics-inspired layer is in line with the well-known and used fact that good inductive biases encoded in the NN architecture lead to an easier training procedure, better generalization capabilities and a lower number of parameters, which in turn reduces overfitting risks Goodfellow et al. (2016); Mehta et al. (2019). The present choice of fully convolutional layers Long et al. (2015) itself serves as an encoding of the translational equivariance symmetry Cohen and Welling (2016) and of the locality of the evolution, allowing for applications to arbitrary domain sizes. In the Supplementary Material, it is shown how omitting the divergence layer degrades the predictive capabilities of the CRNN when extrapolating over long-time evolutions.
@@
Based on our tests, the best set of hyperparameters is the following: the CRNN is composed of two stacked Convolutional Gated Recurrent Unit (GRU) Chung et al. (2014); Shi et al. (2015) blocks, each using 10 channels for hidden states and $3\times 3\times 3$ convolution kernels. A sketch of the overall CRNN architecture is shown in Fig. 1(a). The resulting NN model contains $\approx 2.9\times 10^{4}$ trainable parameters, almost an order of magnitude less than the 2D prototype presented in Ref. Lanzoni et al., 2022, despite the more challenging 3D setting here tackled. Such compression is made possible by the modifications previously discussed. The training loss used is the voxel-wise Mean Squared Error between the predicted and ground truth sequence. Training of the NN parameters is performed using the standard implementation of the Adam optimizer Kingma and Ba (2017) with a learning rate of $10^{-5}$. Before being passed to the CRNN, training sequences are downscaled by a factor of $1/2$ in all spatial dimensions using nearest interpolation to ease GPU memory requirements during training. As it will be shown in the following, the NN prediction accuracy is not significantly impacted by such a downscale process, which, on the other hand, can be leveraged to tackle larger computational cells at reduced computational costs, as has also been demonstrated in previous works Lanzoni et al. (2022); Fan et al. (2024). Data augmentation based on reflections, 90°90°90  ° degree rotations and the $\mathbf{Z}_{2}$ symmetry $\varphi\rightarrow(1-\varphi)$ is also exploited in training. We point again readers interested in implementation details to the GitHub repository, where the full code used to perform training and evolutions here reported is available.
@@
Fig. 1(b) reports training and validation loss respectively. To perform the full training, $\approx 50h$ are required using a workstation using a single Nvidia RTX A4000 GPU. Discrepancies between the two values in the initial training stages are due to the use of a curriculum learning schedule Bengio et al. (2009); Goodfellow et al. (2016), which proved to speed up training convergence. Specifically, at the beginning of training the model parameters are optimized using a simplified loss function, i.e. the CRNN is required to generate only the last state provided the previous 49 ones. In the following 25 epochs, the loss function gets gradually more complex as the CRNN is required to generate progressively longer sequences. At the end of this curriculum stage, the ML model has to reproduce the full sequence based only on the initial condition.
@@
Figure 2: (a) Spatial generalization test. The domain size is 8 times larger than that used in the training set (twice as large in all directions). One-to-one accordance with the ground truth evolution can still be observed. (b) Evolution on a computational cell with the same size of training for a $\varphi$ initial profile shaped as the Stanford bunny. In the upper half of the snapshots, only $\varphi\geq 0.5$ is shown. MSE values for the whole sequences are reported in the insets.